# Numerical study of hypershadows in higher–dimensional black holes

### v2

@@ 5 Conclusion
Extensive studies have clarified the optical appearance of four–dimensional black holes, culminating in direct observational confirmation by the EHT. This naturally raises the question of how such features generalize to higher–dimensional settings. In this work, we introduced a numerical approach based on backward ray tracing to compute the hypershadow, the three–dimensional generalization of the black hole shadow in five–dimensional spacetimes. Our method enables systematic control over observer parameters and facilitates the reconstruction of the full shadow as a three–dimensional volumetric object. The method allows systematic control of observer parameters and reconstructs the hypershadow as a volumetric object. For visualization, we combined discrete sampling with surface contouring and introduced reflection difference maps to quantify mirror symmetries.
Applying this framework to the Schwarzschild–Tangherlini and Myers–Perry solutions, we confirmed the exact spherical symmetry of the Schwarzschild–Tangherlini hypershadow and revealed the contrasting behaviors of Myers–Perry geometries. The cohomogeneity–one case ($a=b$) exhibits rigid rotational invariance, while the single–rotation case ($a=0$) shows clear size reduction, displacement, and symmetry breaking that grow with inclination and spin. To quantify these effects, we introduced two observables: a distortion parameter, measuring size deficit relative to the Tangherlini baseline, and a displacement parameter, characterizing global offsets.
Our findings demonstrate that hypershadows provide a powerful probe of the rotational structure and symmetry properties of higher–dimensional black holes. The numerical framework presented here establishes a robust foundation for future numerical explorations of more exotic objects, such as black rings and their prospective toroidal hypershadows.
Acknowledgements

### v3

@@ 5 CONCLUSION
Extensive studies have clarified the optical appearance of four-dimensional black holes, culminating in direct observational confirmation by the Event Horizon Telescope (EHT). This naturally raises the question of how such features generalize to higher-dimensional settings. In this work, we introduced a numerical approach based on backward ray tracing to compute the hypershadow, the three-dimensional generalization of the black hole shadow in five-dimensional spacetimes. Our method enables systematic control over observer parameters and facilitates the reconstruction of the full shadow as a three-dimensional volumetric object. The method allows systematic control of observer parameters and reconstructs the hypershadow as a volumetric object. For visualization, we combined discrete sampling with surface contouring and introduced reflection difference maps to quantify mirror symmetries.
Applying this framework to the Schwarzschild-Tangherlini and Myers-Perry solutions, we confirmed the exact spherical symmetry of the Schwarzschild-Tangherlini hypershadow and revealed the contrasting behaviors of Myers-Perry geometries. The cohomogeneity-one case ($a=b$) exhibits rigid rotational invariance, while the single-rotation case ($a=0$) shows clear size reduction, displacement, and symmetry breaking that grow with inclination and spin. To quantify these effects, we introduced two observables: a distortion parameter, measuring size deficit relative to the Tangherlini baseline, and a displacement parameter, characterizing global offsets.
Our findings demonstrate that hypershadows provide a powerful probe of the rotational structure and symmetry properties of higher-dimensional black holes. The numerical framework presented here establishes a robust foundation for future numerical explorations of more exotic objects, such as black rings and their prospective toroidal hypershadows.
ACKNOWLEDGMENTS